Exercise 5-9 Compute and Use the Degree of Operating Leverage (L05-8] Engberg Company installs lawn sod in home yards. The coRequired 1 Required 2 Required 3 What is the companys degree of operating leverage? (Round your answer to 1 decimal place.)Required 1 Required 2 Required 3 Using the degree of operating leverage, estimate the impact on net operating income of a 5%Required 1 Required 2 Required 3 Construct a new contribution format income statement for the company assuming a 5% increase
Answer #1

a) Degree of operating leverage = Contribution margin/Net income = 48000/10000 = 4.8

b) Net operating income increase by (4.8*5) = 24%

c) Contribution margin income statement

Amount Percent of sales
Sales 84000 100%
Variable cost 33600 40%
Contribution margin 50400 60%
Fixed cost 38000
Operating income 12400